A 60 g mass with a horizontal velocity of 50 cm/s collides with an 85 g mass at rest. After collision, the 60 g mass travel 30 O below the horizontal and the 85 g mass travels 45 O above the horizontal. Find the speed of each mass after impact.
1
Expert's answer
2020-12-17T04:51:46-0500

So,


X: m1v1=m1v1cos30°+m2v2cos45°m_1v_1=m_1v_1'\cos30°+m_2v_2'\cos45°

Y: m1v1sin30°=m2v2sin45°m_1v_1'\sin30°=m_2v_2'\sin45°


We get


v2=m1v1m2sin45°sin30°+m2sin45°=0.060.50.085sin45°sin30°+0.085sin45°=0.17(m/s)v_2'=\frac{m_1v_1}{\frac{m_2\sin45°}{\sin30°}+m_2\sin45°}=\frac{0.06\cdot 0.5}{\frac{0.085\cdot\sin45°}{\sin30°}+0.085\cdot\sin45°}=0.17(m/s)


v1=m2v2sin45°m1sin30°=0.0850.17sin45°0.06sin30°=0.34(m/s)v_1'=\frac{m_2v_2'\sin45°}{m_1\sin30°}=\frac{0.085\cdot 0.17\cdot\sin45°}{0.06\cdot\sin30°}=0.34(m/s)